OpenStudy (turingtest):

Pinocchio’s prehistory. Master Ciliegia received an order to make a certain number of stools. “If I make three stools a day, starting from today,” the carpenter thought aloud, “I'll finish on Sunday. If I make five stools a day, I'll be done on Friday.”—“And what day is it today?” asked a curious talking block of wood. Indeed, what day is it?

OpenStudy (anonymous):

wednesday

OpenStudy (turingtest):

A little proof? You have a 1/7 chance just guessing, so I think some demonstration is needed.

OpenStudy (anonymous):

 24 - 3x = 30 -5x 2x = 6 x = 3